Billie Bob purchased a used camera (five-year property) for use in his sole proprietorship in the...

  1. Billie Bob purchased a used camera (five-year property) for use in his sole proprietorship in the prior year. The basis of the camera was $2,400. Billie Bob used the camera in his business 60 percent of the time during the first year. During the second year, Billie Bob used the camera 40 percent for business use. Calculate Billie Bob’s depreciation deduction during the second year, assuming the sole proprietorship had a loss during the year. (Billie Bob did not place the asset in service in the last quarter.) (Use MACRS Table
  1. $0
  2. $48
  3. $192
  4. $336
  5. None of the choices are correct

I need a explanation of how to get Year 2 Depreciation Expense. Please provide a detailed breakdown of how to solve the problem please!
Answer #1

As per IRS code, when the business use of property drops below 50%, the straight line method shall be used for the upcoming years. The prior years’ excess depreciation has to be restated i.e. recaptured.

The asset’s recovery period is 5 years.

The asset is not placed in service in last quarter. Hence, half-year convention would apply to the asset.

Current year depreciation without prior year adjustment as per half-year convention for 5-year asset = $ 2400 * 0.20 * 40% = $ 192.

The prior year depreciation needs to be recaptured.

Prior year depreciation to be recaptured = 2400*0.1*60% = $ 144

Hence, the current year depreciation expense would be $ 192-144 = $ 48.

The correct choice is b) $ 48

Note that rate depreciation is as per Table as per 5 year property in Half-Year convention for Year 1. This is Table A-8 where depreciation is straight line and convention used is half-year. The rate is 20% for year 2 and 10% for Year 1 that is to say, current year rate is 20% and prior year rate would be 10%.

kindly upvote